t=0 100V 0.22 } 1F + 0.25H3

Solve for v(t), t>0.
a. Find the initial conditions.
b. Write the differential equation.
c. Find the general form of the solution.
d. Find the coefficients of the solution by matching the initial conditions.
Answer #1

too Iswitch closed), Бл b . Oian § nir) hallol 100V initial condition 10.10) 5 100 = 20 amp : 5 -100+502107+UeCOV ..Melonta 1e tyremat te doen zo 0+20+ de 20 20. 20 3 - A, - 44 2 20 - A144A 2 20 - - A2+4AL - 342=20 A2 = 20 A1=-A2 - 20 Vitia e uk vit)tuts 240.241 Tower 0.25 laswol overdamped system- a>Wo Si= -x + √2² - wo2 Soz - 2.5+ 12.53”-1034 -2.5 - V12.511.1934 1-4 Ar e